When the stress in the plane stress state is given by ?_〖?〗 = 15 MPa, ?_〖?〗 = 75 MPa, and ?_〖??〗 = 40 MPa, how many MPa are the values ​​of the main stresses?
Answer #1

Calculate the main stresses (principle stresses). -0 012 0,70 2 + + Tv 2 15+ 75 2 + 15-75 2 +(40) =(45) + (-30)* +(40) = (45)